What is the single discount equivalent of a 20/15 series discount
Answer #1

The single discount equivalent to 20/15 series discount = 100 - [(80/100)*(85/100)] = 100 - 0.68 = 0.32 or 32%
Answer #2
The original value is 100 According the single discount series = 100-20 =80 And 100-15 = 85 Therefore, the single discount equivalent of 20/15 = 100-(80/100)*(85/100) = 100-0.68 = 0.32 and percentage is 32
answered by: anonymous